# MagicDrive: Street View Generation with Diverse 3D Geometry Control

### v1

@@ 5.2 Main Results
Realism and Controllability Validation. To confirm the efficacy of MagicDrive in generating realistic images, we employ the nuScenes validation set to synthesize street-view images and show the metrics in Table 1. MagicDrive surpasses all other methods in image quality, delivering significantly lower FID scores. Despite the inherent domain gap between stable diffusion pre-training and street views, our refinement process effectively mitigates this, yielding highly convincing results. Regarding controllability, assessed via BEV segmentation tasks, MagicDrive consistently equals or exceeds baseline results at 224$\times$400 resolution. This is mainly attributed to the distinct encoding design which significantly boosts generation precision on vehicles. At 272$\times$736 resolution, advancements in our encoding strategy are demonstrated by enhanced vehicle mIoU performance. Cropping large areas of each view affects the road mIoU on CVT negatively. Nonetheless, BEVFusion’s results in 3D object detection further corroborate the efficacy of our bounding box encoding.
@@ 5.3 Qualitative Evaluation
Comparison with Baselines. We assessed MagicDrive against two baselines, BEVGen and BEVControl, synthesizing multi-camera views for the same validation scenes. Figure 5 illustrates that MagicDrive generates images markedly superior in quality to both baselines, particularly excelling in accurate object positioning and maintaining consistency in street views for backgrounds and objects, notably over BEVControl. Such performance primarily stems from MagicDrive ’s bounding box encoder and its cross-view attention module.
Multi-level Controls. The design of MagicDrive introduces multi-level controls to street-view generation through separation encoding. This section demonstrates the capabilities of MagicDrive by exploring three control signal levels: scene level (time of day and weather), background level (BEV map alterations and conditional views), and foreground level (object orientation and deletion). As illustrated in Figure 1 and Figure 6, MagicDrive adeptly accommodates alterations at each control level, maintaining multi-camera consistency and high realism in generation.
@@ 6 Ablation Study
Bounding Box Encoding. MagicDrive utilizes separate encoders for bounding boxes and road maps. To demonstrate the efficacy, we train a ControlNet (Zhang et al., 2023a) that takes the BEV map with both road and object semantics as a condition (like BEVGen), denoted as “w/o $E_{box}$” in Table 4. Given the relatively small size of objects in the BEV map, it is necessary for a separate $E_{box}$ to effectively help the model accurately represent vehicle annotations, evidenced by the performance gap in vehicle mIoU. Besides, by applying visible object filter $f_{viz}$, both the road and the vehicle mIoU improve significantly. This is due to that visibility filtering reduces the burden of optimization. Additionally, we explored a variant of MagicDrive, incorporating $E_{box}$ alongside BEV containing both road and object semantics. However, this redundancy did not enhance performance, reinforcing the merit of integrating diverse types of information through different strategies.
Effect of Classifier-free Guidance. We focus on the two most crucial conditions, i.e. object boxes and road maps, and analyze how CFG affects the performance of generation. We change CFG from 1.5 to 4.0 and plot the change of validation results from CVT in Figure 7. Firstly, by increasing the CFG scale, FID keeps degrading, since CFG significantly changes the contrast and sharpness, which is also observed by previous works (Chen et al., 2023b). Secondly, by keeping the map the same for both conditional inference and unconditional inference, we can eliminate the effect of CFG on the map condition. As shown by blue lines of Figure 7, increasing CFG scale results in the highest vehicle mIoU at CFG=2.5, but the road mIoU keeps decreasing. Thirdly, with $M=\{0\}$ for unconditional inference in CFG, road mIoU significantly increases. However, it slightly degrades the guidance on vehicle generation. As stated in Section 4.3, CFG gains complication through incorporating more conditions. Although we ease the case for training, we still face different choices of CFG during inferece. We leave the in-depth investigation for this case as future work.
@@ 7 Conclusion
Acknowledgement. This work is supported in part by General Research Fund (GRF) of Hong Kong Research Grants Council (RGC) under Grant No. 14203521 and in part by the Research Matching Grant Scheme under Grant No. 8601109. We gratefully acknowledge the support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.
@@ Appendix B More Experimental Details
Semantic Classes for Generation. To support most perception models on nuScenes, we try to include semantics commonly used in most settings (Huang et al., 2021; Zhou & Krähenbühl, 2022; Liu et al., 2023; Ge et al., 2023). Specifically, for objects, ten categories include car, bus, truck, trailer, motorcycle, bicycle, construction vehicle, pedestrian, barrier, and traffic cone. For the road map, eight categories include drivable area, pedestrian crossing, walkway, stop line, car parking area, road divider, lane divider, and roadblock.
@@ Appendix C Ablation on Number of Attending Views
Appendix D More Experiments with 3D Object Detection
@@
0.5$\times$ w/o synthetic data 29.92 32.62 too few epochs too few epochs
w/ MagicDrive 32.88 (+2.96) 36.49 (+3.87)
1$\times$ w/o synthetic data 32.48 37.61 64.92 69.42
w/ MagicDrive 35.14 (+2.66) 39.63 (+2.02) 67.28 (+2.36) 70.14 (+0.72)
2$\times$ w/o synthetic data 35.16 40.49 67.87 71.12
w/ MagicDrive 35.40 (+0.24) 41.23 (+0.74) 68.15 (+0.28) 71.18 (+0.06)
@@
Appendix E More Discussion
@@
Note MagicDrive-generated street views by now can only perform as augmented samples to train with real data. It is exciting to train perception models totally with generated data, which will be explored in the future. Meanwhile, more flexible usage of the generated street views beyond data augmentation, especially incorporation with generative pre-training (Chen et al., 2023a; Zhili et al., 2023), contrastive learning (Chen et al., 2021; Liu et al., 2022b) and DM-based perception models (Chen et al., 2022; Ji et al., 2023), is also an appealing future research direction.
Appendix F More Generation Results

### v6

@@ 5.2 Main Results
Realism and Controllability Validation. We assess MagicDrive’s capability to create realistic street-view images with the annotations from the nuScenes validation set. As shown by Table 1, MagicDrive outperforms others in image quality, yielding notably lower FID scores. Regarding controllability, assessed via BEV segmentation tasks, MagicDrive equals or exceeds baseline results at 224$\times$400 resolution due to the distinct encoding design that enhances vehicle generation precision. At 272$\times$736 resolution, our encoding strategy advancements enhance vehicle mIoU performance. Cropping large areas negatively impacts road mIoU on CVT. However, our bounding box encoding efficacy is backed by BEVFusion’s results in 3D object detection.
@@ 5.3 Qualitative Evaluation
Comparison with Baselines. We assessed MagicDrive against two baselines, BEVGen and BEVControl, synthesizing multi-camera views for the same validation scenes (the comparison with BEVGen is in the Appendix D). Figure 5 illustrates that MagicDrive generates images markedly superior in quality to BEVControl, particularly excelling in accurate object positioning and maintaining consistency in street views for backgrounds and objects. Such performance primarily stems from MagicDrive ’s bounding box encoder and its cross-view attention module.
Multi-level Controls. The design of MagicDrive introduces multi-level controls to street-view generation through separation encoding. This section demonstrates the capabilities of MagicDrive by exploring three control signal levels: scene level (time of day and weather), background level (BEV map alterations and conditional views), and foreground level (object orientation and deletion). As illustrated in Figure 1, Figure 6, and Appendix E, MagicDrive adeptly accommodates alterations at each level, maintaining multi-camera consistency and high realism in generation.
5.4 Extension to Video Generation
We demonstrate the extensibility of MagicDrive to video generation by fine-tuning it on nuScenes videos. This involves modifying self-attention to ST-Attn (Wu et al., 2023a), adding a temporal attention module to each transformer block (Figure 7 left), and tuning the model on 7-frame clips with only the first and the last frames having bounding boxes. We sample initial noise independently for each frame using the UniPC (Zhao et al., 2023) sampler for 20 steps and illustrate an example in Figure 7 right.
Furthermore, by utilizing the interpolated annotations from ASAP (Wang et al., 2023c) like DriveDreamer (Wang et al., 2023b), MagicDrive can be extended to 16-frame video generation at 12Hz trained on Nvidia V100 GPUs. More results (e.g., video visualization) can be found on our website.
Figure 7: Extend MagicDrive to video generation. Left: changes in transformer block for video generation. Right: result of video generation. Only keyframes have bounding box control.
@@ 6 Ablation Study
Bounding Box Encoding. MagicDrive utilizes separate encoders for bounding boxes and road maps. To demonstrate the efficacy, we train a ControlNet (Zhang et al., 2023a) that takes the BEV map with both road and object semantics as a condition (like BEVGen), denoted as “w/o $E_{box}$” in Table 4. Objects in BEV maps are relatively small, which require separate $E_{box}$ for accurate vehicle annotations, as shown by the vehicle mIoU performance gap. Applying visible object filter $f_{viz}$ significantly improves both road and vehicle mIoU by reducing the optimization burden. A MagicDrive variant incorporating $E_{box}$ with BEV of road and object semantics didn’t enhance performance, emphasizing the importance of integrating diverse information through different strategies.
Effect of Classifier-free Guidance. We focus on the two most crucial conditions, i.e. object boxes and road maps, and analyze how CFG affects the performance of generation. We change CFG from 1.5 to 4.0 and plot the change of validation results from CVT in Figure 8. Firstly, by increasing CFG scale, FID degrades due to notable changes in contrast and sharpness, as seen in previous studies (Chen et al., 2023c). Secondly, retaining the same map for both conditional and unconditional inference eliminates CFG’s effect on the map condition. As shown by blue lines of Figure 8, increasing CFG scale results in the highest vehicle mIoU at CFG=2.5, but the road mIoU keeps decreasing. Thirdly, with $M=\{0\}$ for unconditional inference in CFG, road mIoU significantly increases. However, it slightly degrades the guidance on vehicle generation. As mentioned in Section 4.3, CFG complexity increases with more conditions. Despite simplifying training, various CFG choices exist during inference. We leave the in-depth investigation for this case as future work.
@@ 7 Conclusion
Acknowledgement. This work is supported in part by the General Research Fund (GRF) of Hong Kong Research Grants Council (RGC) under Grant No. 14203521, in part by the CUHK SSFCRS funding No. 3136023, and in part by the Research Matching Grant Scheme under Grant No. 7106937, 8601130, and 8601440. We gratefully acknowledge the support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research. This research has been made possible by funding support from the Research Grants Council of Hong Kong through the Research Impact Fund project R6003-21.
@@ Appendix B More Experimental Details
Semantic Classes for Generation. To support most perception models on nuScenes, we try to include semantics commonly used in most settings (Huang et al., 2021; Zhou & Krähenbühl, 2022; Liu et al., 2023a; Ge et al., 2023). Specifically, for objects, ten categories include car, bus, truck, trailer, motorcycle, bicycle, construction vehicle, pedestrian, barrier, and traffic cone. For the road map, eight categories include drivable area, pedestrian crossing, walkway, stop line, car parking area, road divider, lane divider, and roadblock.
@@ Appendix C Ablation on Number of Attending Views
Appendix D Qualitative Comparison with BEVGen
Figure 12 illustrates that MagicDrive generates images with higher quality compared to BEVGen (Swerdlow et al., 2023), particularly excelling in objects. Such enhancement can be attributed to MagicDrive’s utilization of the diffusion model and the adoption of a customized condition injection strategy.
Figure 12: Qualitative comparison with BEVGen on driving scene from nuScenes validation set. We highlight some areas with rectangles to ease comparison. Compared with BEVGen, image quality of objects from MagicDrive is much better.
Appendix E More results with control from different conditions
Figure 13 shows scene level control (time of day) and background level control (BEV map alterations). MagicDrive can effectively reflect these changes in control conditions through the generated camera views.
Figure 13: Showcase for scene-level control with MagicDrive. The scene is from the nuScenes validation set.
Appendix F More Experiments with 3D Object Detection
@@
0.5$\times$ w/o synthetic data 30.21 32.76 too few epochs too few epochs
w/ MagicDrive 33.29 (+3.08) 36.69 (+3.93)
1$\times$ w/o synthetic data 32.88 37.81 65.40 69.59
w/ MagicDrive 35.40 (+2.52) 39.76 (+1.95) 67.86 (+2.46) 70.72 (+1.13)
2$\times$ w/o synthetic data 35.49 40.66 68.33 71.31
w/ MagicDrive 35.74 (+0.25) 41.40 (+0.74) 68.58 (+0.25) 71.34 (+0.03)
@@
Appendix G More Discussion
@@
Note that MagicDrive-generated street views can currently only perform as augmented samples to train with real data, and it is exciting to train detectors solely with generated data, which will be explored in the future. More flexible usage of the generated street views beyond data augmentation, especially incorporation with generative pre-training (Chen et al., 2023a; Zhili et al., 2023), contrastive learning (Chen et al., 2021; Liu et al., 2022b) and the large language models (LLMs) (Chen et al., 2023b; Gou et al., 2023), is an appealing future research direction. It is also interesting to utilize the geometric controls in different circumstances beyond 3D scenarios (e.g., multi-object tracking (Li et al., 2023a) and concept removal (Liu et al., 2023b)).
Appendix H Detailed Analysis on 3D Object Detection with Synthetic Data
Table 7: Per-class performance comparison with BEVFusion for 3D object detection with 1×1$\times$1 × setting. Results are tested on the nuScenes validation set.
Data mAP car cone barrier bus ped. motor. truck bicycle trailer constr.
BEVFusion 32.88 50.67 50.46 48.62 37.73 35.74 30.40 27.54 24.85 15.56 7.28
$+$ MagicDrive 35.40 51.86 53.56 51.15 40.43 38.10 33.11 29.35 27.85 18.74 9.83
Difference +2.52 +1.20 +3.10 +2.53 +2.70 +2.36 +2.71 +1.81 +3.00 +3.19 +2.55
We provide per-class AP for 3D object detection from the nuScenes validation set using BEVFusion in Table 7. From the results, we observe that, firstly, the improvements for large objects are significant, for example, buses, trailers, and construction vehicles. Secondly, objects with less diverse appearances, such as traffic cones and barriers, show more improvement, especially compared to trucks. Thirdly, we note that the improvement is marginal for cars, while significant for pedestrians, motorcycles, and bicycles. This may be because the baseline already performs well for cars. For pedestrians, motorcycles, and bicycles, even though distant objects from the ego car are generated less faithfully, MagicDrive can synthesize high-quality objects near the ego car, as shown in Figure 17-18. Therefore, more accurate detection of objects near the ego car contributes to improvements for these classes. Overall, mAP improvement comes with promotion in all classes’ AP, indicating MagicDrive can indeed help the training of perception models.
Appendix I More results for BEV segmentation
BEVFusion is also capable of BEV segmentation and considers most of the classes we used in the BEV map condition. Due to the lack of baselines, we present the results in Table 8 to facilitate comparison for future works. As can be seen, the 272$\times$736 resolution does not outperform the 224$\times$400 resolution. This is consistent with the results from CVT in Table 1 on the Road segment. Such results confirm that better map controls rely on maintaining the original aspect ratio for generation training (i.e., avoiding cropping on each side).
Table 8: Generation fidelity to BEV map conditions. Results are tested with BEVFusion for BEV segmentation on the nuScenes validation set.
Methods resolution mIOU for 6 classes
CAM-only CAM+LiDAR
Oracle - 57.09 62.94
Oracle 224$\times$400 52.72 58.49
MagicDrive 224$\times$400 30.24 48.21
MagicDrive 272$\times$736 28.71 47.12
Appendix J Generalization of Camera Parameters
To improve generalization ability, MagicDrive encodes raw camera intrinsic and extrinsic parameters for different perspectives. However, the generalization ability is somewhat limited due to nuScenes fixing camera poses for different scenes. Nevertheless, we attempt to exchange the intrinsic and extrinsic parameters between the three front cameras and three back cameras. The comparison is shown in Figure 14. Since the positions of the nuScenes cameras are not symmetrical from front to back, and the back camera has a ${}^{\circ}$ FOV compared to the ${}^{\circ}$ FOV of the other cameras, clear differences between front and back views can be observed for the same 3D coordinates.
Figure 14: To show the generalization ability of learned camera encoding, we exchange the camera parameters between 3 front cameras and 3 back cameras. The 3D position is the same for two generations. We highlight some areas (with box boxes)
Appendix K More Generation Results